A 300 uF capacitor separated by a distance of 4.0 mm was charged to a potential difference of 250 V. What is the energy density in its region in terms of Joules/m³?

1
Expert's answer
2022-04-28T16:42:49-0400
"\omega=\frac{W}{V}=\frac{CV^2}{2V}=\frac{\epsilon_0V^2}{2d^2}"

"=\frac{8.85*10^{-12}*250^2}{2*0.004^2}=0.017\:\rm J/m^3"